# Broken symmetry solutions in one-dimensional lattice models via many-body perturbation theory

### v1

@@ Abstract
In this work we study self-consistent solutions in one-dimensional lattice models obtained via many-body perturbation theory. The Dyson equation is solved in a fully self-consistent manner via the algorithmic inversion method based on the sum-over-pole representation (AIM-SOP) of dynamical operators. We start by validating our self-consistent AIM-SOP implementation by taking as test case the one-dimensional Hubbard model. We then move to the study of antiferromagnetic and charge density wave solutions in one-dimensional lattice models, taking into account a long-range Coulomb interaction between the electrons. Complementary, by solving the Sham-Schlüter equation, we can compute the non-interacting Green’s function reproducing the same charge density of the interacting system. In turn, this allows for the evaluation of the derivative discontinuity in the Kohn-Sham potential, which gives a measure of how the Kohn-Sham gap approximates the many-body fundamental gap of these systems.
@@ I Introduction
One-dimensional systems constitute an optimal playground to study the effect of electronic correlations at a reduced computational cost with respect to systems of higher dimensionality. A recent study of one-dimensional hydrogen chains Motta et al. (2020) has shown how the presence of long-range Coulomb repulsion between the electrons can give rise to a rich phase diagram, including antiferromagnetic (AFM) domains and metal-to-insulator (MIT) transitions. Furthermore one-dimensional quasi-metals constitute a prototypical example of Peierls transitions Peierls (1929), where a periodic lattice distortion destabilizes the metallic phase leading to an insulating state with a lower total energy.
@@
However it has been recently pointed out that the same systems can host a purely electronic charge density wave (CDW) that does not involve a structural symmetry breaking Barborini et al. (2022). This electronic CDW, which destabilizes the one-dimensional quasi-metal by opening a small energy gap could be related to an excitonic mechanism Sherrington and Kohn (1968); Varsano et al. (2017). However studies performed on one-dimensional Hubbard chains Hubbard (1963) have shown that an approximate treatment of the electron-electron repulsion can lead to unphysical solutions with a broken symmetry ground state Joost et al. (2022); Honet et al. (2022): this is the case of antiferromagnetic (AFM) solutions found via many-body perturbation theory Fetter and Walecka (1971); Martin et al. (2016); Stefanucci and van Leeuwen (2013), while an analytical treatment of one-dimensional Hubbard clusters predicts a non-magnetic (NM) ground state Lieb and Wu (1968, 2003a) (also referred to as a paramagnetic (PM) state in the literature Honet et al. (2022)).
@@
In the framework of MBPT, the information about the charged excitations of the interacting system is encoded in the one-particle Green’s function (GF), with the effect of the electronic interaction described by the non-local and dynamical self-energy operator $\Sigma(\omega)$. For conserving approximations Baym and Kadanoff (1961); Baym (1962), the MBPT problem can also be recast in a functional formulation via the Luttinger-Ward Luttinger and Ward (1960) and Klein Klein (1961) functionals, which ultimately require the Dyson equation to be solved in a self-consistent way. Despite the potential of MBPT, density functional theory (DFT) Hohenberg and Kohn (1964) in its Kohn-Sham formulation (KS-DFT) Kohn and Sham (1965) is still one of the mostly used tools in electronic structure calculations Marzari et al. (2021). The bridge between the MBPT framework and the KS-DFT formalism is given by the Sham-Schlüter equation (SSE) Sham (1985); Sham and Schlüter (1983) that allows one to compute the local Kohn-Sham exchange-correlation potential once the self-consistent GF and its corresponding self-energy are known. Although this approach is still seldom adopted, the SSE is mostly solved in its linearized form Casida (1995).
@@
Recently, Chiarotti et al. Chiarotti et al. (2022); Chiarotti (2023); Chiarotti et al. (2024) have exploited sum-over-poles (SOP), i.e. meromorphic (and rational) functions of the frequency having only simple (first order) poles, to represent dynamical propagators. Similar constructions have also been employed, e.g., in Refs. Engel et al. (1991); Puig von Friesen et al. (2010); Di Sabatino et al. (2021). Moreover, in Refs. Chiarotti et al. (2022); Chiarotti (2023); Chiarotti et al. (2024), the SOP representation is combined with the algorithmic inversion method (AIM-SOP) providing an in-principle exact way of solving the Dyson equation at all frequencies by means of the diagonalization of a linear eigenvalue problem in a larger space Güttel and Tisseur (2017). Similar approaches have been used in the context of DMFT, GW, and Bethe-Salpeter equation Savrasov et al. (2006); Budich et al. (2012); Wang et al. (2012); Bintrim and Berkelbach (2021, 2022). In addition, the AIM-SOP framework can also be exploited Ferretti et al. (2024) to provide a fully analytical way of evaluating the Klein functional Klein (1961).
@@
In this work we study self-consistent solutions of the Dyson equation as resulting from different self-energy approximations in a MBPT context. We apply the framework to one-dimensional lattice models, also including long-range interactions beyond the local Hubbard repulsion. Additionally, we also solve the (non-linearized) SSE and evaluate the derivative discontinuity of the Kohn-Sham potential. Technically, we make use of the AIM-SOP framework in order to implement a fully self-consistent solution of the Dyson equation, to analytically evaluate the $\Phi[G]$ terms of the Klein functional, and to solve the SSE. Our results show that the AIM-SOP framework is capable of solving self-consistently the Dyson equation in a stable and accurate way for the systems considered, obtaining an excellent agreement with Refs. Puig von Friesen et al. (2010); Di Sabatino et al. (2021); Joost et al. (2022). Physically, when a long-range electronic repulsion is introduced in the model, both AFM or CDW solutions can be found depending on the strength of the non-local interaction term, consistently with the results of Refs. Motta et al. (2020); Barborini et al. (2022). By solving the SSE problem, we also find that the KS potential displays a significant derivative discontinuity when CDW solutions are found, in agreement with Ref. Schonhammer and Gunnarsson (1987).
@@
This manuscript is organized as follows: in Sec. II we give an overview of the MBPT formalism. In Sec. III we explain how the Dyson equation is solved in a fully self-consistent way via the AIM-SOP procedure. In Sec. IV we test the accuracy and performance of the AIM-SOP procedure using one-dimensional Hubbard chains as test case and we compare our results against those of other authors. In Sec. V we study the presence of broken symmetry (AFM and CDW) solutions in periodic one-dimensional lattices in presence of a non-local electron-electron repulsion. Additional and complementary details are reported in the Appendixes and in the Supplemental Material sup .
@@ II.1 MBPT Theory
The one-body Green’s function Fetter and Walecka (1971); Stefanucci and van Leeuwen (2013); Martin et al. (2016) is a dynamical operator that allows one to evaluate expectation values of one-body observables, charged excitations (one particle spectrum), and total energies, e.g. through the Galitskii-Migdal (GM) formula Galitskii and Migdal (1958) or the Luttinger-Ward and Klein functionals Luttinger and Ward (1960); Klein (1961). In this work we focus on zero-temperature systems at equilibrium. Starting from a non-relativistic independent-particle Hamiltonian $h_{0}$, with the non-interacting Green’s function $G_{0}$ written as
@@
the effect of the electron-electron interaction term $v$ is taken into account through the introduction of the self-energy $\Sigma(\omega)$, which is in general a non-local and dynamical operator. The GF of the interacting system is then obtained from the Dyson equation
@@
Approximations to the self-energy (SE) operator can be obtained, e.g., within MBPT, where the SE can be diagrammatically defined by considering selected classes of diagrams. In this regard, Baym and Kadanoff Baym (1962); Baym and Kadanoff (1961) have shown that conserving approximations for $\Sigma=\Sigma[G]$ are obtained when one can introduce a functional $\Phi[G]$ such that
@@
Examples of these functionals are given by the Luttinger-Ward Luttinger and Ward (1960) and Klein Klein (1961) functionals. We will focus on the latter, which is given in terms of $G$ by:
@@
$+\text{Tr}_{\omega}\left\{I-G_{0}^{-1}G\right\}+\Phi[G].$
While these functionals assume in general different values Dahlen et al. (2006); Lani and Marzari (2024) for a generic $G$, they all return the ground state total energy at their stationary points. Notably, the $G$ that makes the functional stationary is obtained by solving a self-consistent Dyson equation, where the SE depends on $G$ via Eq. (3). Here and in the following the $\text{Tr}_{\omega}$ symbol denotes the operation of taking the frequency integration of the trace of the operator considered. It is also helpful to stress that the $\Phi$ functional can be built through a non-diagrammatic and non-perturbative procedure Potthoff (2006), thereby being well suited to study broken symmetry solutions of systems in the strong correlation regime. The self-energy approximations considered in this work are the Hartree-Fock (HF), second Born (2B), and GW approximations. Their SOP representation is reported in Appendix A.
@@ II.2 Bridging MBPT and KS-DFT frameworks
As stated in Sec. I, it is possible to bridge MBPT and KS-DFT via the SSE. One defines the KS GF through the use of a local and static potential $v_{KS}(\mathbf{r})$ via
$G_{KS}(\omega)=\left[\omega I-h_{0}-v_{KS}\right]^{-1}.$ (5)
The local potential $v_{KS}(\mathbf{r})$ is determined in such a way to reproduce the same charge density of the interacting system. This is obtained by solving the following Sham-Schlüter equation (SSE) Sham and Schlüter (1983); Sham (1985):
@@
where the dependence of the SE on the self-consistent GF is highlighted. Since the potential $v_{KS}$ is defined up to a constant, corresponding to a rigid energy shift leaving the density unchanged, we require $G_{KS}$ and $G$ to share the same Fermi level.
Furthermore the evaluation of $G_{KS}$ through Eq. (5) allows one to determine the value of the discontinuity $\Delta_{xc}$ in the KS potential for a given SE approximation at the MBPT level Sham and Schlüter (1983); Godby et al. (1986). This space-independent quantity relates the energy gap $E^{KS}_{g}$ of the KS system to the physical fundamental gap, i.e. the quasiparticle energy gap $E^{QP}$ evaluated at the MBPT level
$E^{QP}_{g}-E^{KS}_{g}=\Delta_{xc}=v^{(+)}_{KS}-v^{(-)}_{KS},$ (7)
with $v^{(\pm)}_{KS}$ being the KS potential associated to an infinitesimal increase or decrease in the number of particles.
In passing, we notice that a linearized version of the SSE, where the self-consistent $G$ is replaced by $G_{KS}$ everywhere (also in the self-energy dependence) has been originally proposed by Casida Casida (1995) and it is used in the context of optimized effective potential (OEP) approaches Kümmel and Kronik (2008); Martin et al. (2016); Ferretti et al. (2014); Hellgren et al. (2018); Vacondio et al. (2022). A dynamical generalization of the SSE has also been proposed by Gatti et al. Gatti et al. (2007) and it has been shown to be flexible enough to reproduce the spectral density of interacting systems Gatti et al. (2007); Ferretti et al. (2014); Vanzini et al. (2018, 2020). The SOP expression for the linear system defined by Eq. (II.2) is reported in Appendix B.
@@ III The AIM-SOP approach
III.1 Sum-over-pole and algorithmic inversion
In the context of Green’s function methods, the SOP representation of propagators or dynamical operators is a meromorphic representation, typically involving only simple poles. This has been recently exploited in Ref. Chiarotti et al. (2022) to study the homogeneous electron gas, and also used in combination with localized GW in a first-principles context Chiarotti (2023); Chiarotti et al. (2024). Analogous meromorphic constructions have also been employed elsewhere when dealing with Green’s function methods Cini and Verdozzi (1987); Engel et al. (1991); Puig von Friesen et al. (2010); Di Sabatino et al. (2021). In particular, the work of Engel and Farid Engel et al. (1991) shows how the SOP representation can also be seen as a truncated Lehmann representation. In Refs. Chiarotti et al. (2022); Chiarotti (2023); Chiarotti et al. (2024) the SOP representation is combined with an algorithmic inversion method (AIM-SOP) to solve the Dyson equation exactly and at all frequencies. In the AIM-SOP approach the non-linear eigenvalue problem (NLEP) described by the Dyson equation is recast into a linear eigenvalue problem in a larger space. The same AIM-SOP approach allows one to evaluate analytically the Klein functional and the RPA correlation energy Ferretti et al. (2024). In this section, based on the work of Refs. Chiarotti et al. (2022, 2024); Ferretti et al. (2024), we give a self-contained description of the AIM-SOP framework.
Adopting a lattice representation, which is the focus of this work, with $i$ and $j$ denoting site indices, the interacting GFs and the self-energy are expressed as
@@ III.1 Sum-over-pole and algorithmic inversion
with simple poles $z_{s}$ and $\Omega_{m}$, and residues defined by the matrix elements $A^{s}_{ij}$ and $\Gamma^{m}_{ij}$, respectively. Regarding the GF, the SOP representation then coincides with the Lehmann spectral representation Lehmann (1954) of a dynamical propagator of a finite system with discrete one-particle excitations, for which the singularities of the GF are isolated and do not merge into a branch cut on the real axis. In passing we stress that, at variance with Refs. Chiarotti et al. (2022); Chiarotti (2023); Chiarotti et al. (2024), we only consider propagators having real poles, while the AIM-SOP framework allows one to treat the general case of complex valued poles. With real poles, the time-ordering of propagators can be defined once a Fermi level $\mu$ is specified Fetter and Walecka (1971); Martin et al. (2016); Stefanucci and van Leeuwen (2013); Farid (1999), by setting, e.g. for the Green’s function, $z_{s}=\epsilon_{s}+i\,\text{sgn}(\mu-\epsilon_{s})\eta$, with $\epsilon_{s}\in\mathbb{R}$ and $\eta\rightarrow 0^{+}$.
The advantage of working with a meromorphic representation of dynamical operators is that several quantities of interest can be evaluated analytically. For instance the matrix elements of the one-body reduced density matrix $\gamma_{ij}$ can be determined from
@@
where the sum is restricted over the “occupied” poles of the dynamical operators – i.e. the poles at energies below the Fermi level of their respective operator. Similar expressions are easily obtained for the spectral function $A_{ij}(\omega)$ Fetter and Walecka (1971); Martin et al. (2016); Stefanucci and van Leeuwen (2013) and for the Galitskii-Migdal formula Galitskii and Migdal (1958). In addition, a dynamical convolution of two SOP operators is still a SOP operator Chiarotti et al. (2022) and its structure in terms of poles and residues can be determined analytically. As stated before, and following Refs. Farid (1999); Onida et al. (2002); Chiarotti et al. (2022, 2024), the problem of solving the Dyson equation, Eq. (2), is equivalent to the following NLEP:
@@
where the eigenvalues $z_{s}$ give the poles of the GF, while the eigenvectors $|f_{s}\rangle$ correspond to the Dyson orbitals and the matrix elements of their projectors give the residues of the GF at the corresponding pole $A^{s}_{ij}~{}=~{}\langle i|f_{s}\rangle\langle f_{s}|j\rangle$ (here written in the case of a Hermitian NLEP). For a generic operator $\Sigma(z)$, the eigenvalues defined by the NLEP of Eq. (11) will be in general complex valued Güttel and Tisseur (2017), and left and right eigenvectors need to be introduced.
@@
where, for a physical self-energy, the factorization is always possible due to the positive semi-definiteness (PSD) of the residues $\Gamma^{m}$ Stefanucci and van Leeuwen (2013). The GF of Eq. (12) can then be interpreted as the GF of a system coupled to a non-interacting bath: the poles of the self-energy give the degrees of freedom of the non-interacting bath, while the operators $V_{m}$ and $V^{\dagger}_{m}$ give the couplings between the bath and the one-particle system. The diagonalization of the overall system-plus-bath Hamiltonian returns the poles of the GF as eigenvalues and the Dyson orbitals as the projection of the eigenvectors on the system subspace Chiarotti et al. (2024).
Finally the AIM-SOP framework allows one to evaluate the Klein functional of Eq. (4) in a fully analytical manner Ferretti et al. (2024). In particular, all the expressions involving the $\text{Tr}_{\omega}\text{Ln}$ term can be computed as
@@
where the sum is restricted over the “occupied” poles and each pole is weighted with its corresponding rank $r_{i}$ (i.e., its degeneracy). In general, the expression in Eq. (III.1) can be used to evaluate the $\text{Tr}_{\omega}\text{Ln}$ term involving any two dynamical operators linked by a Dyson equation, as in Eq. (2), with a kernel $\Sigma$ having at most simple poles $\Omega_{m}$. This is used in practice to evaluate the RPA correlation energy of the GW approximation as discussed in Ref. Ferretti et al. (2024). It is relevant to stress that the poles appearing in Eq. (III.1) are weighted with their ranks, and not with their residues: this means that poles with limited weight, such as those from satellites, count as much as quasiparticle poles in the evaluation of the total energy. In this respect, numerical accuracy and consistency in the solution of the Dyson equation are critical when evaluating Eq. (III.1). Having a stable value from the evaluation of Eq. (III.1) can be taken as a sign of having a good resolution in the satellite poles.

### v2

@@ Abstract
In this work we study self-consistent solutions in one-dimensional lattice models obtained via many-body perturbation theory. The Dyson equation is solved in a fully self-consistent manner via the algorithmic-inversion method based on the sum-over-poles representation (AIM-SOP) of dynamical operators. In particular, we focus on the GW approximation, analyzing the spectral properties and the emergence of possible magnetic- or charge-density-wave broken symmetry solutions. We start by validating our self-consistent AIM-SOP implementation by taking as test case the one-dimensional Hubbard model. We then move to the study of antiferromagnetic and charge density wave solutions in one-dimensional lattice models, taking into account a long-range Coulomb interaction between the electrons. We show that moving from local to non-local electronic interactions leads to a competition between antiferromagnetic and charge-density-wave broken symmetry solutions. Complementary, by solving the Sham-Schlüter equation, we can compute the non-interacting Green’s function reproducing the same charge density of the interacting system. In turn, this allows for the evaluation of the derivative discontinuity of the Kohn-Sham (KS) potential, showing that its contribution to the fundamental gap can become dominating in some of the studied cases.
@@ I Introduction
One-dimensional systems constitute an optimal playground to study the effect of electronic correlations at a reduced computational cost with respect to systems of higher dimensionality. A recent study of one-dimensional hydrogen chains [1] has shown how the presence of long-range Coulomb repulsion between the electrons can give rise to a rich phase diagram, including antiferromagnetic (AFM) domains and metal-to-insulator (MIT) transitions. Furthermore one-dimensional quasi-metals constitute a prototypical example of Peierls transitions [2], where a periodic lattice distortion destabilizes the metallic phase leading to an insulating state with a lower total energy.
@@
However it has been recently pointed out that the same systems can host a purely electronic charge density wave (CDW) that does not involve a structural symmetry breaking [3]. This electronic CDW, which destabilizes the one-dimensional quasi-metal by opening a small energy gap could be related to an excitonic mechanism [4, 5]. However studies performed on one-dimensional Hubbard chains [6] have shown that an approximate treatment of the electron-electron repulsion can lead to unphysical solutions with a broken symmetry ground state [7, 8]: this is the case of antiferromagnetic (AFM) solutions found via many-body perturbation theory [9, 10, 11], while an analytical treatment of one-dimensional Hubbard clusters predicts a non-magnetic (NM) ground state [12, 13] (also referred to as a paramagnetic state in the literature [8]).
@@
In the framework of MBPT, the information about the charged excitations of the interacting system is encoded in the one-particle Green’s function (GF), with the effect of the electronic interaction described by the non-local and dynamical self-energy operator $\Sigma(\omega)$. For conserving approximations [14, 15], the MBPT problem can also be recast in a functional formulation via the Luttinger-Ward [16] and Klein [17] functionals, which ultimately require the Dyson equation to be solved in a self-consistent way. Despite the potential of MBPT, density functional theory (DFT) [18] in its Kohn-Sham formulation (KS-DFT) [19] is still one of the mostly used tools in electronic structure calculations [20]. The bridge between the MBPT framework and the KS-DFT formalism is given by the Sham-Schlüter equation (SSE) [21, 22] that allows one to compute the local Kohn-Sham exchange-correlation potential once the self-consistent GF and its corresponding self-energy are known. Although the SSE can be solved self-consistently, the solution is typically found in its linearized form [23].
@@
Part of the difficulty in solving the self-consistent Dyson equations stems from the handling of dynamical potentials. In this respect, Chiarotti et al. [24, 25, 26] have recently exploited sum-over-poles (SOP), i.e. meromorphic (and rational) functions of the frequency having only simple (first order) poles, to represent dynamical propagators and self-energies. Similar constructions have also been employed, e.g., in Refs. [27, 28, 29]. Moreover, in Refs. [24, 25, 26], the SOP representation is combined with the algorithmic-inversion method (AIM-SOP) providing an in-principle exact way of solving the Dyson equation at all frequencies by means of the diagonalization of a linear eigenvalue problem in a larger space [30]. Similar approaches have been used in the context of DMFT, GW, Bethe-Salpeter equation [31, 32, 33, 34, 35, 36]. For instance, SOP-representations have been recently exploited in the multi-pole approximation (MPA) for full-frequency GW calculations [37, 38, 39] to represent the RPA response functions, self-energies, and Green’s functions. In addition, the AIM-SOP framework can also be exploited [40] to provide a fully analytical way of evaluating the Klein functional [17].
@@
In this work we study self-consistent solutions of the Dyson equation as resulting from different self-energy approximations in a MBPT context, applied to one-dimensional lattice models, also including long-range interactions beyond the local Hubbard repulsion. Additionally, we also solve the – non-linearized – SSE and evaluate the derivative discontinuity of the Kohn-Sham potential. Technically, we make use of the AIM-SOP framework [24, 26, 25] in order to implement a fully self-consistent solution of the Dyson equation, and also exploit a recently developed analytically expression to evaluate the $\Phi[G]$ terms of the Klein functional [40]. Our results show that the AIM-SOP framework is capable of solving self-consistently the Dyson equation in a stable and accurate way for the systems considered, obtaining an excellent agreement with Refs. [28, 29, 7].
Physically, when a long-range electronic repulsion is introduced in the model, both AFM and CDW solutions can be found depending on the strength of the non-local interaction term, consistently with the results of Refs. [1, 3]. We show that the interplay between the local and non-local electronic interactions leads to a competition between AFM and CDW phases in the exact ground state, and thus to an explicit symmetry breaking in the case of MBPT solutions. This effect is particularly relevant in the case of a non-interacting metallic ground state. Remarkably, the evaluation of the total energy within the GW approximation improves considerably in the case of symmetry unrestricted calculations. Furthermore, by solving the SSE problem, we also find that the KS potential displays a significant derivative discontinuity when CDW solutions are found, in agreement with Ref. [41], stressing the role of exchange-correlation corrections to the fundamental gap of the KS system.
@@
This manuscript is organized as follows: in Sec. II we give an overview of the MBPT formalism. In Sec. III we explain how the Dyson equation is solved in a fully self-consistent way via the AIM-SOP procedure. In Sec. IV we test the accuracy and performance of the AIM-SOP procedure using one-dimensional Hubbard chains as test case and we compare our results against those of other authors. In Sec. V we study the presence of broken symmetry (AFM and CDW) solutions in periodic one-dimensional lattices in presence of a non-local electron-electron repulsion. Additional and complementary details are reported in the Appendixes and in the Supplemental Material [42] (see also references [7, 43, 40, 44, 45, 46] therein).
@@ II.1 MBPT Theory
The one-body Green’s function [9, 11, 10] is a dynamical operator that allows one to evaluate expectation values of one-body observables, charged excitations (one particle spectrum), and total energies, e.g. through the Galitskii-Migdal (GM) formula [45] or the Luttinger-Ward and Klein functionals [16, 17]. In this work we focus on zero-temperature systems at equilibrium. Starting from a non-relativistic independent-particle Hamiltonian $h_{0}$, with the non-interacting Green’s function $G_{0}$ written as
@@
the effect of the electron-electron interactions is taken into account through the introduction of the self-energy $\Sigma(\omega)$, which is in general a non-local and dynamical operator. The GF of the interacting system is then obtained from the Dyson equation
@@
Approximations to the self-energy (SE) operator can be obtained, e.g., within MBPT, where the SE can be diagrammatically defined by considering selected classes of diagrams. In this regard, Baym and Kadanoff [15, 14] have shown that conserving approximations for $\Sigma=\Sigma[G]$ are obtained when one can introduce a functional $\Phi[G]$ such that
@@
The $\Phi[G]$ term can then be used within the Luttinger-Ward [16] or Klein [17] functionals to express the total energy of the system. We will focus on the latter, which is given in terms of $G$ by:
@@
$+\text{Tr}_{\omega}\left\{I-G_{0}^{-1}G\right\}+\Phi[G],$
where we use the notation $_{\omega}[\dots]=\int\frac{d\omega}{2\pi i}e^{i\omega 0^{+}}\text{Tr}[\dots]$ Tr [ … ]. While these functionals assume in general different values [47, 48] for a generic $G$, they both return the ground state total energy at their stationary points. Notably, the $G$ that makes the functional stationary is obtained by solving a self-consistent Dyson equation, where the SE depends on $G$ via Eq. (3). It is also helpful to stress that the $\Phi$ functional can be built through a non-diagrammatic and non-perturbative procedure [49], thereby being well suited to study broken symmetry solutions of systems in the strong correlation regime. The self-energy approximations considered in this work are the Hartree-Fock (HF), second Born (2B), and GW approximations. Their SOP representation is reported in Appendix A.
@@ II.2 Bridging MBPT and KS-DFT frameworks
As stated in Sec. I, it is possible to bridge MBPT and KS-DFT via the SSE. One defines the KS GF through the use of a local and static potential $v^{\text{KS}}(\mathbf{r})$ via
$G^{\text{KS}}(\omega)=\left[\omega I-h_{0}-v^{\text{KS}}\right]^{-1}.$ (5)
The local potential $v^{\text{KS}}(\mathbf{r})$ is determined in such a way to reproduce the same charge density of the interacting system. This is obtained by solving the following Sham-Schlüter equation (SSE) [22, 21]:
@@
where the dependence of the SE on the self-consistent GF is highlighted. Since the potential $v^{\text{KS}}$ is defined up to a constant, corresponding to a rigid energy shift leaving the density unchanged, we require $G^{\text{KS}}$ and $G$ to share the same Fermi level.
Furthermore the evaluation of $G^{\text{KS}}$ through Eq. (5) allows one to determine the value of the discontinuity $\Delta_{xc}$ in the KS potential for a given SE approximation at the MBPT level [22, 50, 51]. This space-independent quantity relates the energy gap $E^{\text{KS}}_{g}$ of the KS system to the physical fundamental gap, i.e. the quasiparticle energy gap $E^{\text{QP}}$ evaluated at the MBPT level
$E^{\text{QP}}_{g}-E^{\text{KS}}_{g}=\Delta_{xc}=v^{\text{KS}}_{+}-v_{-}^{\text{KS}},$ (7)
with $v_{\pm}^{\text{KS}}$ being the KS potentials associated to an infinitesimal increase or decrease in the number of particles.
In passing, we notice that a linearized version of the SSE, where the self-consistent $G$ is replaced by $G^{\text{KS}}$ everywhere (also in the self-energy dependence) has been originally proposed by Casida [23] and it is used in the context of optimized effective potential (OEP) approaches [52, 10, 53, 54, 55]. A dynamical generalization of the SSE has also been proposed by Gatti et al. [56] and it has been shown to be flexible enough to reproduce the spectral density of interacting systems [56, 53, 57, 58]. The SOP expression for the linear system defined by Eq. (II.2) is reported in Appendix B.
@@ III The AIM-SOP approach
III.1 Sum-over-pole and algorithmic-inversion method
In the context of Green’s function methods, the SOP representation of propagators or dynamical operators is a meromorphic representation, typically involving only simple poles. This has been recently exploited in Ref. [24] to study the homogeneous electron gas, and also used in a first-principles context in combination with a localized GW formulation [25, 26]. Analogous meromorphic constructions have also been employed elsewhere when dealing with Green’s function methods [59, 27, 28, 29]. In particular, the work of Engel and Farid [27] shows how the SOP representation can also be seen as a truncated Lehmann representation. In Refs. [24, 25, 26] the SOP representation is combined with an algorithmic-inversion method (AIM-SOP) to solve the Dyson equation exactly and at all frequencies. In the AIM-SOP approach the non-linear eigenvalue problem (NLEP) described by the Dyson equation is recast into a linear eigenvalue problem in a larger space. The same AIM-SOP approach allows one to evaluate analytically the Klein functional and the RPA correlation energy [40]. In this section, based on the work of Refs. [24, 26, 40], we give a self-contained description of the AIM-SOP framework.
Adopting a lattice representation, which is the focus of this work, with $i$ and $j$ denoting site indices, the interacting GF and the self-energy are expressed as
@@ III.1 Sum-over-pole and algorithmic-inversion method
with simple poles $z_{s}$ and $\Omega_{m}$, and residues defined by the matrix elements $A^{s}_{ij}$ and $\Gamma^{m}_{ij}$, respectively. Regarding the GF, the SOP representation then coincides with the Lehmann representation [60] of a dynamical propagator of a finite system with discrete one-particle excitations, for which the singularities of the GF are isolated and do not merge into a branch cut on the real axis. In passing we stress that, at variance with Refs. [24, 25, 26], we only consider propagators having real poles, while the AIM-SOP framework allows one to treat the general case of complex valued poles. With real poles, the time-ordering of propagators can be defined once a Fermi level $\mu$ is specified [9, 10, 11, 61], by setting, e.g. for the Green’s function, $z_{s}=\epsilon_{s}+i\,\text{sgn}(\mu-\epsilon_{s})\eta$, with $\epsilon_{s}\in\mathbb{R}$ and $\eta\rightarrow 0^{+}$.
The advantage of working with a meromorphic representation of dynamical operators is that several quantities of interest can be evaluated analytically. For instance the matrix elements of the one-body reduced density matrix $\gamma_{ij}$ can be determined as
@@
where the sum is restricted to the “occupied” poles of the dynamical operators – i.e. the poles at energies below the Fermi level. Similar expressions are easily obtained for the spectral function $A_{ij}(\omega)$ [9, 10, 11] and for the Galitskii-Migdal formula [45]. In addition, a dynamical convolution of two SOP operators is still a SOP operator [24] and its structure in terms of poles and residues can be determined analytically. As stated before, and following Refs. [61, 62, 24, 26], the problem of solving the Dyson equation, Eq. (2), is equivalent to the following non-linear eigenvalue problem (NLEP):
@@
where the eigenvalues $z_{s}$ give the poles of the GF, while the eigenvectors $|f_{s}\rangle$ correspond to the Dyson orbitals and the matrix elements of their projectors give the residues of the GF at the corresponding pole, $A^{s}_{ij}~{}=~{}\langle i|f_{s}\rangle\langle f_{s}|j\rangle$ (here written in the case of a Hermitian NLEP). For a generic operator $\Sigma(z)$, the eigenvalues defined by the NLEP of Eq. (11) will be in general complex valued [30], and left and right eigenvectors need to be introduced.
@@
where, for a physical self-energy, the above factorization is always possible due to the positive semi-definiteness (PSD) of the residues $\Gamma^{m}$ [11]. The GF of Eq. (12) can then be interpreted as the GF of a system coupled to a non-interacting bath: the poles of the self-energy give the degrees of freedom of the non-interacting bath, while the operators $V_{m}$ and $V^{\dagger}_{m}$ provide the coupling between the bath and the one-particle system. The diagonalization of the overall system-plus-bath Hamiltonian returns the poles of the GF as eigenvalues and the Dyson orbitals as the projection of the eigenvectors on the system subspace [26].
Finally the AIM-SOP framework allows one to evaluate the Klein functional of Eq. (4) in a fully analytical manner [40]. In particular, all the expressions involving the $\text{Tr}_{\omega}\text{Ln}$ term can be computed as
@@
where the sum is restricted over the “occupied” poles and each pole is weighted with its corresponding rank $r_{i}$ (i.e., its degeneracy). In general, the expression in Eq. (III.1) can be used to evaluate the $\text{Tr}_{\omega}\text{Ln}$ term involving any two dynamical operators linked by a Dyson equation, as in Eq. (2), with a kernel $\Sigma$ having at most simple poles $\Omega_{m}$. This is used in practice to evaluate the RPA correlation energy of the GW approximation as discussed in Ref. [40]. It is relevant to stress that the poles appearing in Eq. (III.1) are weighted with their ranks, and not with their residues: this means that poles with limited weight, such as those from satellites, count as much as quasiparticle poles in the evaluation of the total energy. In this respect, numerical accuracy and consistency in the solution of the Dyson equation are critical when evaluating Eq. (III.1). Moreover, having a stable value from the evaluation of Eq. (III.1) can be taken as a sign of having a good resolution in the satellite poles.